# Refined Policy Distillation: From VLA Generalists to RL Experts

### v1

@@ Abstract
Recent generalist Vision-Language-Action Models (VLAs) can perform a variety of tasks on real robots with remarkable generalization capabilities. However, reported success rates are often not on par with those of expert policies. Moreover, VLAs usually do not work out of the box and often must be fine-tuned as they are sensitive to setup changes. In this work, we present Refined Policy Distillation (RPD), an RL-based policy refinement method that enables the distillation of large generalist models into small, high-performing expert policies. The student policy is guided during the RL exploration by actions of a teacher VLA for increased sample efficiency and faster convergence. Different from previous work that focuses on applying VLAs to real-world experiments, we create fine-tuned versions of Octo and OpenVLA for ManiSkill2 to evaluate RPD in simulation. As our results for different manipulation tasks demonstrate, RPD enables the RL agent to learn expert policies that surpass the teacher’s performance in both dense and sparse reward settings. Our approach is even robust to changes in the camera perspective and can generalize to task variations that the underlying VLA cannot solve.
@@ I Introduction
Recently, robot learning has seen a paradigm shift from Reinforcement Learning (RL) of task-specific policies to large generalist Vision-Language-Action Models (VLAs) that are trained in a supervised fashion using imitation learning techniques [1, 2, 3, 4, 5, 6, 7, 8]. VLAs are easy to train and show good performance on different real-world manipulation tasks with a single model. However, they usually need to be trained on a large number of human expert demonstrations, limiting their performance to the amount and variety of the training data. Different from the field of natural language processing, creating more data is very cost-intensive in robotics as it usually means collecting trajectories by humans through time-consuming teleoperation. Due to limited training data, VLAs have problems generalizing across setups [9]. Applying VLAs in new environments, thus, often requires fine-tuning on domain-specific data, requiring human effort yet again to record demonstrations, which further limits their widespread adaptation. Only recently, Xu et al. [10] utilized data generated by RL agents for VLA fine-tuning, which can significantly scale up setup-specific fine-tuning.
@@
Figure 1: The architecture of RPD: We distill a VLA policy into an RL student policy. This has two effects: First, the RL agent is bootstrapped with guided exploration through the VLA teacher. Second, the student policy can interact with the environment and is, thus, able to surpass the VLA’s performance, refining the distilled policy.
In this work, we introduce Refined Policy Distillation (RPD) as a novel method for distilling and optimizing expert policies from generalist VLA policies. The idea behind RPD is that a VLA policy trained on large-scale datasets stores general task knowledge that can inform the training process of an RL algorithm, even if the policy by itself can solve the task only at a low or possibly even zero success rate. Our evaluation shows that RPD is not only capable of distilling and improving policies from fine-tuned VLA policies but that it can also successfully distill VLA knowledge while learning unseen tasks. This also applies when the camera angle changes, a situation in which current VLAs largely fail. RPD extends the Proximal Policy Optimization (PPO) [15] algorithm and only requires a slight modification of its loss function.
The main contributions of this work can be summarized as follows:
@@
    We propose RPD, a novel method for distilling and refining large VLAs with on-policy RL into compact expert policies.
@@
    We demonstrate RPD’s increased sample efficiency and improved stability on six different in-distribution tasks for both dense and sparse reward settings.
@@
    We show its positive generalization on two out-of-distribution task variations and a changed camera perspective.
@@ II Related Work
In recent years, several approaches have been developed for learning generalist robot policies conditioned on vision and language inputs. One of the most common ways is to use an existing VLM backbone trained on large amounts of internet-scale data and fine-tune it on recorded robot trajectories [3, 5, 7, 8], most notably from the Open X-Embodiment dataset [16]. Other models are based on large transformers and trained on vision and language or goal-image instructions from scratch [1, 2, 4, 17, 18]. We refer to both of these transformer-based policy types as VLAs.
@@
Behavior Cloning (BC) is an umbrella term for offline imitation learning: Given a dataset of state transitions, the goal is to learn a policy that follows the trajectories from the dataset as closely as possible [24]. Exact formulations differ, but the main idea is to maximize the likelihood of the state transitions in the dataset. BC was also combined with RL to speed up the agent’s exploration through a fixed pre-recorded dataset [25, 26]. Recently, the term Offline RL has been used to describe similar approaches but they focus mainly on off-policy RL methods [27, 28]. Offline RL is related to PD as both approaches aim to learn from expert knowledge but uses datasets instead of learned policies. We employ methods similar to BC to process external actions in the RL training loop but leverage pre-trained generalist policies to generate them.
Xu et al. [10] investigated dataset generation with an offline soft actor-critic variant [28, 29] for VLA fine-tuning. While their experiments show performance improvements of the VLA policy of up to 40%, the method still requires already trained RL expert policies. By contrast, our method distills large foundation models into smaller on-policy RL agents, which avoids training RL policies from scratch. Through interaction with the environment, these policies can surpass the performance of their VLA teachers on specific tasks and become expert policies.
@@ III Methodology
RPD is a method for distilling large generalist VLA policies into small and fast task-specific expert policies. Standard supervised BC techniques usually learn policies that do not exceed the performance of the teacher For this reason, we combine policy distillation with RL. This helps the student agent to explore more efficiently and reach higher rewards faster, even in sparse reward settings. In addition, it also allows for refinement of the teacher policy to surpass its performance and become an expert policy for the distilled task. At the core of RPD is a PPO RL agent with a modified objective function that includes a Mean Squared Error (MSE) term between the action mean predicted by PPO and the expectation of the VLA action to pull the action mean of the RL policy closer to the predictions of the VLA. Fig. 1 provides an overview of how RPD integrates VLA actions into the agent-environment RL training loop.
@@ III-A Refined Policy Distillation
RPD combines RL with a Behavior Cloning (BC) objective. For RL, we model the interaction between an agent and its environment as a finite Markov Decision Process (MDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},r,p,\rho_{0},\gamma)$ [30]. $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ the reward function,$p(s_{t+1}\mid s_{t},a_{t})$ the state transition probability distribution of the environment, $\rho_{0}$ the probability distribution over the initial states $s_{0}$, and $\gamma\in[0,1]$ the discount factor. Every task has its own MDP associated with it. We consider stochastic policies $\pi_{\theta}(a_{t}\mid s_{t})$ that are parameterized by weights $\theta$. The goal of RL is to find a set of weights $\theta^{*}$ that maximizes the expected sum of discounted rewards:
@@
BC trains a policy from expert demonstrations that are usually provided as a dataset $\rho_{\mathcal{D}}$ of state transitions. We adopt a BC objective [25] that maximizes the likelihood of the policy :
@@
Note that the PPO objective is maximized while the MSE is minimized, which is why it appears with a negative sign. In practice, we found that sampling a single VLA action per step already yields good performance. Increasing the sample size would slow down the training considerably and, therefore, was not further investigated. Our training objective is to find a set of optimal weights $\theta^{*}$ that maximize the objective function $\mathcal{L}_{\text{RPD}}$:
@@ III-B Tasks, Dataset, and VLA Fine-Tuning
TABLE I: Evaluation rewards (RW) and success rates (SR) of the VLAs on the selected ManiSkill2 tasks.
@@
We use eight different manipulation tasks from ManiSkill2 [31] to evaluate RPD. Fig. 2 provides a visual overview. Our criteria for choosing these tasks were the availability of demonstrations that are no longer than the maximum episode length for reasons detailed below and that all action spaces contain relative task space actions $a$ with the following format:
@@
The camera perspectives shown in Fig. 2 match the environment state observed by both the VLA and the RL agent. Moreover, the same camera was used to record the VLA fine-tuning data. ManiSkill2 calls this the human camera. It is normally not meant to be used for training as it contains visual hints such as a green sphere for the goal location of the cube in the PickCube task.
@@
Goal locations in ManiSkill2 are drawn randomly and provided to the agent only through explicit state information. The VLAs considered in this work, however, only receive RGB camera state information and a fairly simple natural language instruction. Therefore, at least one of these two input modalities needs to provide a hint for the currently selected target location. While encoding the target in the language instruction appears to be a natural choice, the precision may be limited as VLAs are usually pre-trained on simple language instructions that do not contain exact spatial information. We leverage the visual cues in ManiSkill2’s human camera to encode all relevant task information. Another reason for using this camera is that its perspective shows the scene from the side rather than from the front. This side view corresponds better to the camera perspective used in many setups of the Open X-Embodiment dataset.
@@
Figure 2: Overview of the eight different tasks from ManiSkill2 that were used to distill expert policies with RPD. See Tab. I for the task names (tasks are depicted in row-major order).
Initially, we evaluated the published checkpoints for OpenVLA and Octo on the selected ManiSkill2 tasks using the human camera and the language instructions shown in Tab. I. However, as reported in the table, both models had a success rate of zero on all tested tasks. This real-to-sim gap is expected, as both Octo and OpenVLA were only trained on real-world datasets. When using ManiSkill2’s raytracing renderer, the success rate of Octo increased to more than $2\%$ on PushCube, indicating that some positive domain transfer is possible for more realistic-looking input data. Still, those success rates are too small for a teacher policy, and ManiSkill2 does not support GPU rendering of the human camera with raytracing, which is required for RL.
@@
In order to evaluate RPD under realistic conditions, we first fine-tuned Octo and OpenVLA on the simulated environment using RL-generated expert demonstrations provided by ManiSkill2. All demonstrations were rendered with the human camera at an image resolution of $256\times 256$. During training, we found that increasing the maximum episode length to 300 steps led to considerably higher success rates as the ManiSkill2 dataset contains a subset with episodes longer than the maximum of the 50 steps used for most environments. As long episodes can make RL algorithms behave less stable and also require a higher number of VLA inference steps, we excluded them from fine-tuning, which led to the expected success rates for both Octo and OpenVLA. The results and language instructions for all tasks and models are summarized in Tab. I.
@@
The two tasks PullCubeTool and PokeCube are excluded from the fine-tuning dataset to test cross-task generalization. While the normalized rewards are low, as expected, the reported success rates seem unreasonably high. This is because in these two tasks, the agent is supposed to grasp a tool and use it to either pull or poke a cube. We found that the VLAs were just pulling or pushing the cube to the target location without utilizing the tool, which is counted as a success by the sparse reward implemented by ManiSkill2 as it just compares the distance towards the goal location and not whether the tool was actually used. The dense reward, however, takes tool use into account and accurately shows the decreased performance. Nevertheless, we still decided to keep both tasks. Even if the learned behavior is just pushing and pulling, which the success rate can still accurately display, it still differs from PushCube and PullCube because of distracting objects in the scene. The fact that the language instruction is partly ignored by the VLA in that case is not relevant for testing the generalization behavior of RPD.
@@ III-C VLA Integration Details
To integrate different VLAs into the RL training loop, we defined a generic policy interface that takes a batch of visual observations and language instructions as input and returns a relative task space motion command including the gripper state. A technically challenging part is the software dependency management for the implementations of the different VLAs, as they are partly incompatible with each other and with the dependencies of the RL training code and the simulator. For example, Octo requires a specific version of JAX, which depends on a CUDA version that does not work with the CUDA version that OpenVLA’s PyTorch version needs.
To solve this issue, we created separate Python environments, each of which contains the correct dependencies. The generic VLA interface is exposed via an HTTP API server that runs with the correct environment. During training, the RL agent connects to this server and sends batches of observations.
A new problem that comes with this approach is serialization. Communication can quickly become a bottleneck when large batches of raw image data have to be serialized. If we run both VLA inference and the RL training on the same machine, we use shared memory to avoid serialization.
Lastly, OpenVLA does not support batch processing in its inference code. This limitation further slows down the already large model, as it complicates the implementation of parallel inference. Moreover, the CPU-based preprocessing occupies excessive CPU resources, necessitating manual code optimizations to move it to the GPU. Even after utilizing a parallelization factor of eight (running two OpenVLA instances on each GPU within a four Nvidia H100 node setup), it remains considerably slower than Octo operating on a single Nvidia A40 GPU. Due to these constraints, we were only able to conduct a subset of our experiments using OpenVLA.
@@ IV Experimental Results
We evaluated RPD on two VLAs, namely Octo [4] and OpenVLA [7]. The latter model with its 7B parameters is much larger than the former with only 93M parameters and therefore requires also considerably more computing time, which is exacerbated by the lack of support for batched inference. This is why we provide evaluation results for OpenVLA only for a subset of the tasks. henever the base model is not explicitly mentioned we refer to Octo.
We split the evaluation into four parts: First, we test RPD on the challenging PickCube task to explore the performance differences between loss variants and baselines. We then evaluate RPD and its PPO baseline on a total of six different tasks that are included in the VLA fine-tuning dataset. RL often struggles with sparse rewards due to inefficient exploration. We argue that RPD guides this exploration process and, thus, should also improve sample efficiency and performance on sparse rewards. To test this hypothesis we also run the tasks mentioned above in a sparse reward setting. Finally, we evaluate RPD’s cross-task and cross-setup generalization capabilities by testing it on two tasks that it was not trained on and a variant of the PushCube task with an altered camera perspective. The latter simulates small setup changes that can easily happen in reality and have a considerable effect on the performance of current VLMs.

### v2

@@ Abstract
Vision-Language-Action Models (VLAs) have demonstrated remarkable generalization capabilities in real-world experiments. However, their success rates are often not on par with expert policies, and they require fine-tuning when the setup changes. In this work, we introduce Refined Policy Distillation (RPD), a novel Reinforcement Learning (RL)-based policy refinement method that bridges this performance gap through a combination of on-policy RL with behavioral cloning. The core idea of RPD is to distill and refine VLAs into compact, high-performing expert policies by guiding the student policy during RL exploration using the actions of a teacher VLA, resulting in increased sample efficiency and faster convergence. We complement our method by fine-tuned versions of Octo and OpenVLA for ManiSkill3 to evaluate RPD in simulation. While this is a key requirement for applying RL, it also yields new insights beyond existing studies on VLA performance in real-world settings. Our experimental results across various manipulation tasks show that RPD enables the RL student to learn expert policies that outperform the VLA teacher in both dense and sparse reward settings, while also achieving faster convergence than the RL baseline. Our approach is even robust to changes in camera perspective and can generalize to task variations that the underlying VLA cannot solve. Our code, dataset, VLA checkpoints, and videos are available at https://refined-policy-distillation.github.io
@@ I Introduction
Recently, robot learning has seen a paradigm shift from Reinforcement Learning (RL) of task-specific policies to large generalist Vision-Language-Action Models (VLAs) that are trained in a supervised fashion using imitation learning techniques [1, 2, 3, 4, 5, 6, 7, 8]. VLAs are easy to train and show good performance on different real-world manipulation tasks with a single model. However, they usually need to be trained on a large number of human expert demonstrations, limiting their performance to the amount and variety of the training data. Different from the field of natural language processing, creating more data is very cost-intensive in robotics, as it usually means collecting trajectories by humans through time-consuming teleoperation. Due to limited training data, VLAs have problems generalizing across setups [9]. Applying VLAs in new environments, thus, often involves fine-tuning on domain-specific data, requiring human effort yet again to record demonstrations, which further limits their widespread adoption. Only recently, Xu et al. [10] utilized data generated by RL agents for VLA fine-tuning, which can significantly speed up the adaptation to specific setups.
@@
Figure 1: The architecture of RPD: We distill a VLA teacher into an RL student policy. This has two effects: First, the RL student agent is bootstrapped with guided exploration through the VLA teacher. Second, the RL student policy can interact with the environment and is, thus, able to surpass the VLA’s performance, refining the distilled policy.
In this work, we introduce Refined Policy Distillation (RPD), a novel method for distilling and optimizing expert policies from VLAs by combining on-policy RL with behavioral cloning. The idea behind RPD is that a VLA trained on large-scale datasets stores general task knowledge that can inform the training process of an RL algorithm, even if the policy by itself can solve the task only at a low or possibly even zero success rate. Our evaluation shows that RPD is not only capable of distilling and improving expert policies from fine-tuned VLAs but that it can also successfully distill VLA knowledge while learning unseen tasks. This also applies when the camera angle changes, a situation in which current VLAs largely fail. RPD extends the Proximal Policy Optimization (PPO) algorithm [15] and only requires a small modification of its loss function.
In summary, we make the following contributions:
@@
    We introduce RPD, a novel method for distilling and refining compact task-specific expert policies from large VLAs with on-policy RL.
@@
    We demonstrate RPD’s increased sample efficiency and stability on six different VLA in-distribution tasks for both dense and sparse reward settings.
@@
    We demonstrate generalization to two out-of-distribution task variations and a changed camera perspective.
  4. 4.
    We provide insight into how current VLAs perform in simulation by evaluating RPD with Octo [4] and OpenVLA [7] variants fine-tuned for ManiSkill3.
@@ II Related Work
In recent years, several approaches for learning generalist robot policies conditioned on vision and language inputs have been developed. One of the most common ways is to use an existing VLM backbone trained on internet-scale data and fine-tune it on recorded robot trajectories [3, 5, 7, 8], most notably from the Open X-Embodiment dataset [16]. Other models are based on large transformers and trained on vision and language or goal-image instructions from scratch [1, 2, 4, 17, 18]. We refer to both of these transformer-based policy types as VLAs.
@@
Behavioral Cloning (BC) is an umbrella term for offline imitation learning: Given a dataset of state transitions, the goal is to learn a policy that follows the trajectories from the dataset as closely as possible [24]. Exact formulations differ, but the main idea is to maximize the likelihood of the state transitions in the dataset. BC was also combined with RL to speed up the agent’s exploration through a fixed pre-recorded dataset [25, 26]. Recently, the term Offline RL has been used to describe similar approaches but they focus mainly on off-policy RL methods [27, 28]. Offline RL is related to PD as both approaches aim to learn from expert knowledge but uses datasets instead of learned policies. We employ methods similar to BC to process external actions in the RL training loop but leverage pre-trained generalist policies to generate them.
Xu et al. [10] investigated dataset generation with an offline soft actor-critic variant [28, 29] for VLA fine-tuning. While their experiments show performance improvements of the VLA policy of up to 40%, the method still requires already trained RL expert policies. By contrast, our method distills large foundation models into smaller on-policy RL agents, which avoids training RL policies from scratch. Through interaction with the environment, these policies can surpass the performance of their VLA teachers on their specific tasks and become expert policies.
@@ III Methodology
RPD is a method for distilling tasks from large generalist VLA policies into small and fast task-specific expert policies. Standard supervised BC techniques usually learn policies that do not exceed the performance of the teacher. For this reason, we combine policy distillation with RL. This helps the student agent to explore more efficiently and reach higher rewards faster, even in sparse reward settings. In addition, it also allows for refinement of the teacher policy to surpass its performance and become an expert policy for the distilled task. At the core of RPD is a PPO RL agent with a modified objective function that includes a Mean Squared Error (MSE) term between the action mean predicted by PPO and the expectation of the VLA action to pull the action mean of the RL policy closer to the predictions of the VLA. Fig. 1 provides an overview of how RPD integrates VLA actions into the agent-environment RL training loop.
@@ III-A Refined Policy Distillation
[…]
@@
BC trains a policy from expert demonstrations that are usually provided as a dataset $\rho_{\mathcal{D}}$ of state transitions. We adopt a BC objective [25] that maximizes the likelihood of the policy under $\rho_{\mathcal{D}}$:
@@
Note that the PPO objective is maximized while the MSE is minimized, which is why it appears with a negative sign. In practice, we found that sampling a single VLA action per step to estimate the expectation already yields good performance. Increasing the sample size would slow down the training considerably and, therefore, was not further investigated. Our training objective is to find a set of optimal weights $\theta^{*}$ that maximize the objective function $\mathcal{L}_{\text{RPD}}$:
@@ III-B Tasks, Dataset, and VLA Fine-Tuning
TABLE I: Evaluation rewards (RW) and success rates (SR) of the VLAs on the selected ManiSkill3 tasks.
@@
We use eight different manipulation tasks from ManiSkill3 [31] to evaluate RPD. Fig. 2 provides a visual overview. Our criteria for choosing these tasks were the availability of demonstrations that are no longer than the maximum episode length and that all action spaces contain relative task space actions $a$:
@@
Since the effectiveness of guidance depends on the VLA teacher’s performance, the student may not always benefit during training. In order to be able to appropriately assess the speedup enabled by RPD, we therefore focus our evaluation on tasks that are solvable by vanilla PPO.
@@
The camera perspectives shown in Fig. 2 match the environment state observed by both the VLA and the RL agent. Moreover, the same camera was used to record the VLA fine-tuning data. ManiSkill3 calls this the human camera. It is normally not meant to be used for training as it contains visual hints such as a green sphere for the goal location of the cube in the PickCube task.
@@
Goal locations in ManiSkill3 are drawn randomly and provided to the agent only through explicit state information. The VLAs considered in this work, however, only receive RGB camera state information and a fairly simple natural language instruction. Therefore, at least one of these two input modalities needs to provide a hint for the currently selected target location. While encoding the target in the language instruction appears to be a natural choice, the precision may be limited as VLAs are usually pre-trained on simple language instructions that do not contain exact spatial information. We leverage the visual cues in ManiSkill3’s human camera to encode all relevant task information. Another reason for using this camera is that its perspective shows the scene from the side rather than from the front. This side view corresponds better to the camera perspective used in many setups of the Open X-Embodiment dataset.
@@
Figure 2: Overview of the eight different tasks from ManiSkill3 that were used to distill expert policies with RPD. See Tab. I for the task names (tasks are depicted in row-major order).
[…]
@@
In order to evaluate RPD under realistic conditions, we first fine-tuned Octo and OpenVLA on the simulated environment using RL-generated expert demonstrations provided by ManiSkill3. It is important to note that this step is only required because we evaluate RPD in an environment these models were not trained on. All demonstrations were rendered with the human camera at an image resolution of $\times 256$ 256. During training, we found that increasing the maximum episode length to 300 steps led to considerably higher success rates as the ManiSkill3 dataset contains a subset with episodes longer than the maximum of the 50 steps used for most environments. As long episodes can make RL algorithms behave less stable and also require a higher number of VLA inference steps, we excluded them from fine-tuning, which led to the expected success rates for both Octo and OpenVLA. The results and language instructions for all tasks and models are summarized in Tab. I.
@@
The two tasks PullCubeTool and PokeCube are excluded from the fine-tuning dataset to test cross-task generalization. While the normalized rewards are low, as expected, the reported success rates seem unreasonably high. This is because in these two tasks, the agent is supposed to grasp a tool and use it to either pull or poke a cube. We found that the VLAs were just pulling or pushing the cube to the target location without utilizing the tool. This is counted as a success by the sparse reward implemented in ManiSkill3 which just compares the distance towards the goal location and not whether the tool was actually used. The dense reward, however, takes tool use into account and accurately shows the decreased performance. Nevertheless, we still decided to keep both tasks. Even if the learned behavior is just pushing and pulling, which the success rate can capture, it still differs from PushCube and PullCube because of distracting objects in the scene. The fact that the language instruction is partly ignored by the VLA in that case is not relevant for testing RPD’s generalization performance.
@@ III-C VLA Integration Details
To integrate different VLAs into the RL training loop, we defined a generic policy interface that takes a batch of visual observations and language instructions as input and returns a relative task space motion command including the gripper state. To manage the dependencies of the different VLAs, we created separate Python environments and designed a generic VLA interface that is exposed via an HTTP API server. During training, the RL agent connects to this server and sends batches of observations. A new problem that comes with this approach is serialization. Communication can quickly become a bottleneck when large batches of raw image data have to be serialized. If we run both VLA inference and the RL training on the same machine, we use shared memory to avoid serialization.
Lastly, OpenVLA does not support batch processing in its inference code. This limitation further slows down the already large model, as it complicates the implementation of parallel inference. Moreover, the CPU-based preprocessing consumes substantial computional resources, requiring manual code optimization to offload it to the GPU. Even after utilizing a parallelization factor of eight (running two OpenVLA instances on each GPU within a four Nvidia H100 node setup), it remains considerably slower than Octo operating on a single Nvidia A40 GPU. Due to these constraints, we were only able to conduct a subset of our experiments using OpenVLA.
@@ IV Experimental Results
We evaluated RPD on two VLAs, namely Octo [4] and OpenVLA [7]. The latter model, with its 7B parameters, is much larger than the former, with only 93M parameters, and therefore requires considerably more computing time, which is exacerbated by the lack of support for batched inference. For this reason, we report evaluation results for OpenVLA on only a subset of the tasks. Whenever the base model is not explicitly mentioned, we refer to Octo.
We split the evaluation into four parts: First, we test RPD on the challenging PickCube task to explore the performance differences between loss variants and baselines. We then evaluate RPD and its PPO baseline on a total of six different tasks that are included in the VLA fine-tuning dataset. RL often struggles with sparse rewards due to inefficient exploration. We argue that RPD guides this exploration process and, thus, should also improve sample efficiency and performance on sparse rewards. To test this hypothesis, we also run the tasks mentioned above in a sparse reward setting. Finally, we evaluate RPD’s cross-task and cross-setup generalization capabilities by testing it on two tasks that it was not trained on and a variant of the PushCube task with an altered camera perspective. The latter simulates small setup changes that can easily happen in reality and have a considerable effect on the performance of current VLMs.